# Preparing Lessons for Progressive Training on Language Models

### 17 Jan 2024 v1

@@ Introduction
Transformer Architecture. Here, we introduce the structure of the Transformer, each block of which mainly contains two basic types of layers, i.e., the multi-head self-attention (MHSA) layer, and the feed-forward neural network (FFN) layer. Assuming inputs of the $l$-th layer are query $\mathbf{Q}_{l}\in\mathbb{R}^{d}$, key $\mathbf{K}_{l}\in\mathbb{R}^{d}$, and value $\mathbf{V}_{l}\in\mathbb{R}^{d}$, where $d$ is the dimension, the $l$-th layer can be formulated as
@@
The parameters of the MHSA layer of the $l$-th layer are denoted by $\mathbf{W}_{l}^{\{Q,K,V,O\}}\in\mathbb{R}^{d\times d}$. The weights of the FFN layer are $\mathbf{W}_{l}^{IN}\in\mathbb{R}^{d\times\alpha d}$ and $\mathbf{W}_{l}^{OUT}\in\mathbb{R}^{\alpha d\times d}$, where $\alpha$ is an expanding ratio that is often set to 4. Note that, we neglect biases in the formulation for simplification, which does not affect the generality of the proposed method. Following the above formulation, all weights of the $l$-th layer can be denoted as $\theta_{l}=\mathbf{W}_{l}^{\{Q,K,V,O,IN,OUT\}}$.
@@
Method Probability Density Function
LVPS $P_{\text{LVPS}}(L^{(t)})=\frac{b}{(L^{(t)}+k)^{2}}$
US $P_{\text{US}}(L^{(t)})=\frac{1}{L-N^{(s)}}$
FS $P_{\text{FS}}(L)=1$
@@
In this section, we conduct experiments to validate the performance of our proposed method.
@@
We implement a series of experiments on BERT and GPT for validation. We use AdamW as the optimizer with a learning rate of $10^{-4}$ and weight decay of $10^{-2}$ in all the experiments. We chose the training batch sizes of 768 and 512 for BERT (Devlin et al. 2019) and GPT (Radford et al. 2019) models, respectively. We use the Scratch model, StackBERT, bert2BERT, and LiGO as baselines. Layer numbers of Apollo are [1, 3, 6, 12] and change at epoch [2, 4, 10]. LiGO is warmly trained for 100 steps as claimed in the original paper (Wang et al. 2023b). The training dataset is a concatenation of English Wikipedia and Toronto Book Corpus (Zhu et al. 2015).

### 18 Jan 2024 v2

@@ Introduction
Transformer Architecture. Here, we introduce the structure of the Transformer (Vaswani et al. 2017), each block of which mainly contains two basic layers, (1) the multi-head self-attention (MHSA) layer, (2) the feed-forward neural network (FFN) layer. Assuming inputs of $l$-th layer are query $\mathbf{Q}_{l}\in\mathbb{R}^{d}$, key $\mathbf{K}_{l}\in\mathbb{R}^{d}$, and value $\mathbf{V}_{l}\in\mathbb{R}^{d}$, where $d$ is the dimension, the $l$-th layer can be formulated as
@@
The MHSA and FFN layers are defined as follows:
(1) MHSA Layer. The MHSA layer is another form of the Self-Attention (SA) layer that allows the model to weigh the importance of different words in a sentence, enabling the ability to capture long-range dependencies and contextual information effectively. Instead of using a single attention mechanism, MHSA divides SA into a multi-head structure. Each head attends to different aspects or dependencies in the input sequence. This allows the model to capture different patterns and relationships between words, potentially improving the model’s ability to learn complex dependencies and relationships. Specifically, the parameters of SA in $l$-the layer are $\mathbf{W}_{l}^{\{Q,K,V,O\}}\in\mathbb{R}^{d\times d}$. Then, MHSA separates these parameters into $M$ heads: $\{\mathbf{W}_{l}^{\{Q,K,V,O\},m}\}_{m=1}^{M}$. As a result, we can denote MHSA as
$\operatorname{Att}_{m}(\mathbf{Q}_{l},\mathbf{K}_{l},\mathbf{V}_{l})=$
$~{}~{}~{}\operatorname{softmax}\left(\frac{\mathbf{Q}_{l}\mathbf{W}_{l}^{Q,m}(\mathbf{K}_{l}\mathbf{W}_{l}^{K,m})^{T}}{\sqrt{d_{k}}}\right)\mathbf{V}_{l}\mathbf{W}_{l}^{V,m}\mathbf{W}_{l}^{{O,m}^{T}},$
$\operatorname{MHSA}((\mathbf{Q}_{l},\mathbf{K}_{l},\mathbf{V}_{l}))=\sum^{M}_{m=1}{\operatorname{Att}_{m}(\mathbf{Q}_{l},\mathbf{K}_{l},\mathbf{V}_{l})},$ (4)
where $d_{k}$ is the dimensionality of the key vectors. The attention heads are computed multiple times in parallel with parameters $\mathbf{W}_{l}^{Q,m}$, $\mathbf{W}_{l}^{K,m}$, and $\mathbf{W}_{l}^{V,m}$. These heads will be transformed linearly through a learnable weight matrix $\mathbf{W}_{l}^{O,m}$. At last, MHSA derives the final result by concatenating the head outputs.
@@
(2) FFN Layer. The FFN layer applies a non-linear transformation to the output of the self-attention layer by increasing the dimension and then shrinking to the original size. $\mathbf{X}_{l}\in\mathbb{R}^{d}$ is the input of the $l$-th layer. The weights are $\mathbf{W}_{l}^{IN}\in\mathbb{R}^{d\times\alpha d}$ and $\mathbf{W}_{l}^{OUT}\in\mathbb{R}^{\alpha d\times d}$, where $\alpha$ is an expanding ratio that is often set to 4. The FFN layer can be formed as
$\operatorname{FFN}(\mathbf{X}_{l})=\operatorname{GeLU}(\mathbf{X}_{l}\mathbf{W}_{l}^{IN})\mathbf{W}_{l}^{OUT},$ (5)
where $\operatorname{GeLU}(\cdot)$ is a non-linear activation function (Hendrycks and Gimpel 2016). The FFN layer plays a crucial role in enhancing the ability to model complex patterns, making it a powerful tool for various natural language processing tasks.
Finally, Applying residual connections to both the MHSA and FFN layers is achieved by incorporating addition and layer normalization operations. These crucial steps serve the purpose of safeguarding the model against collapse and overfitting of the training data. Note that, we neglect biases in the formulation for simplification, which does not affect the generality of the proposed method. Following the above formulation, all weights of the $l$-th layer can be denoted as $\theta_{l}=\mathbf{W}_{l}^{\{Q,K,V,O,IN,OUT\}}$.
@@
Method Probability Density Function
LVPS $P_{\text{LVPS}}(L^{(t)})=\frac{b}{(L^{(t)}+k)^{2}}$
US $P_{\text{US}}(L^{(t)})=\frac{1}{L-N^{(s)}}$
FS $P_{\text{FS}}(L)=1$
@@
In this section, we conduct a series of experiments to validate the performance of our proposed method.
@@
We implement a series of experiments on BERT and GPT for validation. We use AdamW as the optimizer with a learning rate of $10^{-4}$ and weight decay of $10^{-2}$ in all the experiments. We chose the training batch sizes of 768 and 512 for BERT (Devlin et al. 2019) and GPT (Radford et al. 2019) models, respectively. We use the Scratch model, StackBERT, bert2BERT, and LiGO as baselines. Layer numbers of Apollo are [1, 3, 6, 12] and change at epoch [2, 4, 10]. LiGO is warmly trained for 100 steps as claimed in the original paper (Wang et al. 2023b). The training dataset is a concatenation of English Wikipedia and Toronto Book Corpus (Zhu et al. 2015). The structures of the BERT and GPT models are in Table 5 of the Appendix.
@@ Acknowledgments
Appendix A Appendix
Hyper-parameter Choice of Edge Sampling (ES)
Edge Sampling (ES) mainly samples low and high depths, while seldom sampling the middle part. As formulated in Table 1, ES can be determined by selecting $k$ as $b$ can be solved when $k$ is chosen. We illustrate the influence of the hyper-parameter $k$ of ES in Fig. 9. We choose $k=10$ for ES in all our experiments, since it offers a clearer differentiation between edges and middle regions, without exhibiting excessive bias towards edges. Hence, it presents a balanced and practical option for comparison.
Figure 9: A case of choosing hyper-parameters $k$ of ES to sample 1-6 layer number.
Structures of Language Models
We show structures of the used language models including the BERT and GPT models in Table 5.
Config BERT-Small BERT-Base BERT-Large GPT-Small GPT-Base
# layers 12 12 24 12 12
# hidden 512 768 1024 512 768
# heads 8 12 16 8 12
# vocab 30522 30522 30522 50257 50257
seq. length 512 512 512 1024 1024
Table 5: The structures of BERT and GPT.